Clark Industries has a defined benefit pension plan that specifies annual, year-end retirement benefits equal to:

1.2% × Service years × Final year’s salary

Stanley Mills was hired by Clark at the beginning of 2002. Mills is expected to retire at the end of 2046 after 45 years of service. His retirement is expected to span 15 years. At the end of 2021, 20 years after being hired, his salary is $87,000. The company’s actuary projects Mills’s salary to be $340,000 at retirement. The actuary’s discount rate is 6%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
1. Estimate the amount of Stanley Mills’s annual retirement payments for the 15 retirement years earned as of the end of 2021.
2. Suppose Clark’s pension plan permits a lump-sum payment at retirement in lieu of annuity payments. Determine the lump-sum equivalent as the present value as of the retirement date of annuity payments during the retirement period.
3. What is the company’s projected benefit obligation at the end of 2021 with respect to Stanley Mills?
4. Even though pension accounting centers on the PBO calculation, the ABO still must be disclosed in the pension disclosure note. What is the company’s accumulated benefit obligation at the end of 2021 with respect to Stanley Mills?
5. If we assume no estimates change in the meantime, what is the company’s projected benefit obligation at the end of 2022 with respect to Stanley Mills?
6. What portion of the 2022 increase in the PBO is attributable to 2022 service (the service cost component of pension expense) and to accrued interest (the interest cost component of pension expense)?

(For all requirements, round final answers to the nearest whole dollars.)
Answer #1

1.)Estimate the amount of Stanley Mills’s annual retirement payments for the 15 retirement years earned as of the end of 2021.

1.2% × Service years × Final year’s salary
1.2% * 20 * 340,000 = 81,600/-

2. Suppose Clark’s pension plan permits a lump-sum payment at retirement in lieu of annuity payments. Determine the lump-sum equivalent as the present value as of the retirement date of annuity payments during the retirement period.

81,600*9.712248988 = 792519/- [ 1- (1+r)^n/r] [1-(1.06)^15/0.06]

3. What is the company’s projected benefit obligation at the end of 2021 with respect to Stanley Mills?

792519 * 0.232998631 = 184656. [ (1.06)^(45-20) = 0.232998631]

4. Even though pension accounting centres on the PBO calculation, the ABO still must be disclosed in the pension disclosure note. What is the company’s accumulated benefit obligation at the end of 2021 with respect to Stanley Mills?

1.2% * 20 * 87,000 = 20,880/-
20,880 * 9.712248988 = 202791.75/-
202792 * 0.232998631 = 47250/-

5. If we assume no estimates change in the meantime, what is the company’s projected benefit obligation at the end of 2022 with respect to Stanley Mills?

1.2% * 21 * 340,000 = 85680
85680*9.7122 = 929263
929263*0.246978 = 229435

6. What portion of the 2022 increase in the PBO is attributable to 2022 service (the service cost component of pension expense) and to accrued interest (the interest cost component of pension expense)?

PBO at the End of 2022 = 229435
PBO at the End of 2021 = -184656
Change in PBO = 44779
Interest Cost = -11079 (184656*6%)

Service Cost = 33700